OpenStudy (anonymous):

integrate: ʃ [e^-(1/2-t)x] / -(1/2-t)^2 dx

OpenStudy (mr.math):

\(\Large\int \frac{e^{-(\frac{1}{2}-t)}x}{-(\frac{1}{2}-t)^2}dx\)?

OpenStudy (anonymous):

yes

OpenStudy (mr.math):

You're integrating with respect to \(x\), you should treat all other variables as constants. That means you can apply the power rule since \(\large \frac{e^{-(\frac{1}{2}-t)}}{-(\frac{1}{2}-t)^2}\) is just like any constant here.

OpenStudy (anonymous):

so what is the answer?

OpenStudy (mr.math):

\(\Large\int \frac{e^{-(\frac{1}{2}-t)}x}{-(\frac{1}{2}-t)^2}dx=\frac{e^{-(\frac{1}{2}-t)}}{-(\frac{1}{2}-t)^2}\cdot \frac{x^2}{2}+c\).

OpenStudy (anonymous):

thats wrong

OpenStudy (mr.math):

It can't be wrong unless the integral is not what I typed. Is \(-(\frac{1}{2}-t)x\) the exponent?

OpenStudy (anonymous):

yes

OpenStudy (mr.math):

That makes a difference.

OpenStudy (mr.math):

\(\Large\int \frac{e^{-(\frac{1}{2}-t)}x}{-(\frac{1}{2}-t)^2}dx=-(\frac{1}{2}-t)\cdot \frac{e^{-(\frac{1}{2}-t)x}}{-(\frac{1}{2}-t)^2}+c\). Now simplify.

OpenStudy (anonymous):

thats wrong as well.

OpenStudy (anonymous):